Mahmoud runs to a marker and back in 15 minutes. His speed on the way to the market is 5m/s and his speed on the way back is 4m/s. Find the distance to the marker.
Answer by mananth(16946)

let distance to market be x m
Total time = 15 minutes= 900 seconds
Time going to market + time returnibg = total time
t=d/r
x/4+x/5 =900
Lcd = 20
Multiply equation by 20
5x+4x= 900*20
9x= 900*20
x= 900*20/9
x= 2000 meters